# Interaction-Aware Model Predictive Decision-Making for Socially-Compliant Autonomous Driving in Mixed Urban Traffic Scenarios

### v1

@@ Abstract
This paper presents the experimental validation of an interaction-aware model predictive decision-making (IAMPDM) approach in the course of a simulator study. The proposed IAMPDM uses a model of the pedestrian, which simultaneously predicts their future trajectories and characterizes the interaction between the pedestrian and the automated vehicle. The main benefit of the proposed concept and the experiment is that the interaction between the pedestrian and the socially compliant autonomous vehicle leads to smoother traffic. Furthermore, the experiment features a novel human-in-the-decision-loop aspect, meaning that the test subjects have no expected behavior or defined sequence of their actions, better imitating real traffic scenarios. Results show that intention-aware decision-making algorithms are more effective in realistic conditions and contribute to smoother traffic flow than state-of-the-art solutions. Furthermore, the findings emphasize the crucial impact of intention-aware decision-making on autonomous vehicle performance in urban areas and the need for further research.
@@ 1 Introduction
Highly automated and autonomous vehicles (AVs) are increasingly becoming a part of our daily lives (Elliott et al., 2019)(Zhou et al., 2022). Integrating these systems into society largely depends on the trust of vulnerable road users, such as cyclists and pedestrians, whose safety is essential (Bouzidi and Hashemi, 2023). Incidents involving automated driving functions often make headlines, fostering public skepticism towards these technologies. Consequently, significant research efforts are directed at enhancing automated vehicles with advanced communication channels and decision-making algorithms designed to manage complex scenarios. For example, in urban environments and city centers – where vehicles move at slower speeds, and pedestrians may cross unexpectedly – an effective human-machine interaction is critical to fostering trust in these systems.
@@
Figure 1: The pedestrian crosses the street at an unsignalised intersection interacting with an autonomous vehicle. With the courtesy of version1 GmbH.
In recent years, advanced algorithms have enabled automated vehicles to interact with pedestrians and cyclists (Hu et al., 2018), (Razmi Rad et al., 2020). The efficiency of these algorithms from the literature is often analyzed either in pure simulation setups (Kampitakis et al., 2023) or in human-in-the-loop tests see e. g. Brand et al. (2024). These human-in-the-loop experiments [^1] Note that the terms experiment and study are used interchangeably in the course of this paper. are usually conducted in virtual environments (Razmi Rad et al., 2020) (Luu et al., 2022). However, literature-based experiments restrict decision possibilities: the vehicle’s behaviors or the test subjects’ expected reactions are predefined, making the experiment more controlled and analysis more accessible, limiting their application in real-world scenarios.
Therefore, this work validates an interaction-aware model predictive decision-making (IAMPDM) within a simulator experiment using a human-in-the-decision-loop setup, where test subjects’ reactions are unspecified to replicate real-world scenarios better. Since field studies with actual vehicles can be risky, human-in-the-loop simulations offer a safer alternative, see e.g Razmi Rad et al. (2020), Tran et al. (2021), Bindschädel et al. (2021) or Kalantari et al. (2023). This paper provides a human-in-the-decision-loop simulator study for the first time with the following contributions:
@@
    Adapting the IAMPDM for the real-time experiments,
@@
    Designing a suitable experimental human-in-the-decision-loop simulator study, and
@@
    Conducting the experiment with 25 participants, offering analysis and recommendations for the real-world implementation of the proposed IAMPDM.
@@ 2 State of the Art
This section provides a short comprehensive literature review of interaction-aware decision-making algorithms, which became large in recent years, since the topic is getting more and more relevant. Therefore, the following state-of-the-art overview focuses on
  • 1.
    decision-making methods, models and algorithms [^2] For simplicity, the terms decision-making methods, models, and algorithms will collectively be referred to as decision-making concepts. and
  • 2.
    experiments and studies, which can verify and validate these concepts.
For more detailed reviews, the reader is referred to Predhumeau et al. (2023).
@@ 2.1 Decision-Making Concepts
In the literature, decision-making concepts can be categorized into two main groups: model-based and model-free approaches.
2.1.1 Model-free Approaches
@@
Model-free concepts do not utilize physics-based or logic-based models. Instead, they learn optimal behaviors through trial and error by directly interacting with the environment. These methods use environmental feedback to refine decision-making, often employing algorithms such as Q-learning or policy gradient methods. Generally, model-free approaches are well-suited for complex scenarios, such as modeling human behavior. A review on this topic can be found in Di and Shi (2021).
@@
In Russo et al. (2021), a Reinforcement Learning (RL) approach is presented addressing pedestrian collision avoidance in autonomous driving systems. The focus is on managing unexpected pedestrian crossings and tracking a specific trajectory. The Deep Deterministic Policy Gradient (DDPG) algorithm is utilized to learn continuous actions. The paper discusses the reward function, agent architecture, and environment model used for training and testing the DDPG-based agent through numerical simulations. Similar, data-based concepts are presented in Raghuram Kadali et al. (2014), Rezaee et al. (2021), Trumpp et al. (2022) and Zhang et al. (2022).
@@
However, the main drawbacks of these model-free concepts are that they cannot easily ensure safety or guarantee the reliability needed in high-risk environments, such as AV-pedestrian interactions. This limitation means that pure learning-based systems are not mature enough to be safely used in real traffic situations. It is to be expected that model-free decision-making approaches will become easier to validate and verify for traffic admission, for which one typical method is the combination of model-free and model-based elements as discussed in Korbmacher and Tordeux (2022) or Derajić et al. (2024).
2.1.2 Model-based Approaches
On the other hand, model-based concepts utilize mathematical models to describe AV-pedestrian interactions using logical or physical principles. These methods rely on this predefined model to predict future states and make decisions. Furthermore, they can handle constraints and uncertainties in the system-design phase, leading to improved safety and reliability. More detailed overviews of model-based methods can be found in Camara et al. (2021), Gulzar et al. (2021) or Predhumeau et al. (2023).
@@ 2.1.2 Model-based Approaches
In Domeyer et al. (2022), a perceptual model is established based on the geometrical relations between drivers and pedestrians, emphasizing the implications for vehicle automation. It explores the coupling of driver and pedestrian. However, only a simulation analysis is provided, with no strong indications for real-world usage. Game theoretical models are presented in Cheng et al. (2015), Skugor et al. (2020) and Pavelko et al. (2022). In these works, the interaction between the pedestrian and vehicle is modeled as a game with two players who optimize their own objective functions by their joint actions. These models can model the interaction between the two players, but they are not suitable to predict their future joint actions.
@@
On the other hand, in Chen et al. (2023), a model predictive control (MPC) formulation is proposed, which can predict the future behavior of the pedestrian. In Škugor et al. (2023), the model predictive formulation is extended with a stochastic component, providing a more realistic overall behavior of the proposed algorithm.
However, these concepts often have a high computational complexity, therefore they are often not suitable for real-time application on a state-of-the-art automotive hardware.
2.2 Validation of Decision-Making Concepts
Since the main contribution of the work is the experimental study with a human-in-the-decision-loop character, the aftercoming part of the overview focuses on pedestrian-autonomous vehicle experiments from the literature. Reviews on the virtual reality studies of pedestrian-automated vehicle interaction can be found in Tran et al. (2021) Zhou et al. (2022).
The two main focuses in the literature are the impact of the external human-machine interface (eHMI) on trust and safety. For instance, in Prattico et al. (2021), various eHMI concepts are compared, and the influence of the eHMI on the test subjects’ trust is investigated. Similarly, in Feng et al. (2023), the impact of eHMI on pedestrian behavior at road crossings is analyzed in a virtual reality setting. The variation of the communication strategies of the automated vehicle is analyzed in Löcken et al. (2019). The authors proposed five groups for the different communication categories, which are implemented and tested in a virtual reality setup. The results showed that the higher the complexity, the more the test subjects trusted the technical system.
The study in Yang et al. (2024) focuses on how pedestrians’ head movements should be interpreted in the presence of automated vehicles at a virtual crossroad, offering insights into non-verbal cues and communication patterns in pedestrian-vehicle interaction for enhanced safety. However, the test subjects in this study are not in the decision loop, which means that the interaction between humans and automation has a negligible effect on the study results.
Thus, the works from the state of the art provide a wide range of experiments and investigations, however, none of these focus on the effect of various decision-making algorithms on the preference of the pedestrians.
@@ 2.3 Shortcoming of the State-of-the-Art Methods
The decision-making and control algorithms from the state of the art have shortcomings that 1) they are either too complex for real-time implementation on general automotive hardware or 2) they are hard to validate in order to pass motor vehicle registration of AVs.
On the other hand, experimental studies from the literature usually use predefined strategies for autonomous vehicles in the case of human-in-the-loop investigations. However, human-in-the-decision-making-loop is an inevitable extension for the further analysis of human factors. This means that the experiments should have no predefined behavior, which can have an impact on the results of the decision-making.
Therefore, our work addresses these challenges and provides an IAMPDM that can run in real time. Furthermore, this work also provides an investigation having a human-in-the-decision-making-loop character.
3 Adjustment of the Model Predictive Algorithm
@@ 3.1 Model Description
The IAMPDM implements an explicit motion model for the pedestrians’ motion. This is necessary to make the decision-making algorithm of the AV more suitable for city-center traffic scenarios with low driving speeds. The core idea of this IAMPDM is presented in Varga et al. (2023b) that we used for our implementation; and we validated the model using the data from Yang et al. (2019).
The proposed model uses a sigmoid function based on the time-to-collision (TTC) values to model the probability of the pedestrian’s velocity changes. In the following, a detailed and more computationally efficient formulation is given. For the explicit motion model, it is assumed that the pedestrian’s choice of speed at the next time step can be modeled by
$\dot{y}_{\mathrm{ped}}(t)=\frac{1}{1+\exp\left(-TTC(t)+c\right)}\cdot v_{\mathrm{ped}}^{\mathrm{ref}},$ (1)
where c is a parameter that can be used to adjust the pedestrian’s characteristics, such as cautious or aggressive behavior. The $TTC$ is the time to collision and is computed by
$TTC(t)=\frac{x_{\mathrm{ped}}(t)-x_{\mathrm{veh}}(t)}{v_{\mathrm{veh}}(t)}-\frac{y_{\mathrm{veh}}(t)-y_{\mathrm{ped}}(t)}{v_{\mathrm{ped}}^{\mathrm{ref}}}.$ (2)
@@
The validity of (1) can be supported by the research of Kadali and Vedagiri (2020) and Forde and Daniel (2021), which propose also velocity-based models for pedestrian-vehicle interactions. The main benefit of the proposed model is that the decision layer is integrated into to control layer, thus the MPC can solve the trajectory planning and decision making tasks in one step.
Gap distance model (see e.g. (Tian et al., 2024)) inherently assumes that the pedestrian’s decision-making is based on the current gap value and assessing the safety without considering how that gap might change in the future. On the other hand, (1) modulates the interaction dynamics between pedestrian and vehicle, which evolves over time. It can account for variables such as speed, acceleration, and changing distances in a continuous manner and used for prediction interaction dynamics.
Assuming a linear dynamics of the vehicle [^3] This assumption is commonly utilized by the application of autonomous vehicles, see e.g. Lynch and Park (2017, Chapter 13). and that the pedestrian moves in the $y$ direction, the following discrete dynamic system is obtained:
@@
$+\underbrace{\begin{bmatrix}0.5\cdot{\Delta}t^{2}\\ \Delta t\\ 0\\ 0\end{bmatrix}}_{{\mathbf{B}}}u_{\mathrm{veh}}(t)+\underbrace{\begin{bmatrix}0\\ 0\\ 0\\ \frac{v_{\mathrm{ped}}^{\mathrm{ref}}}{1+e^{-TTC(t)+c}}\end{bmatrix}}_{\boldsymbol{z}(t)}$ (3)
where the desired acceleration of the vehicle ${u_{\mathrm{veh}}(t)=a_{\mathrm{des}}}$ is the system input. With the model (3.1), the prediction of the future states for $N$ steps is possible using the batch formulation Borrelli et al. (2017):
$\mathcal{A}=\begin{bmatrix}\mathbf{A}\\ \mathbf{A}^{2}\\ \vdots\\ \mathbf{A}^{N}\end{bmatrix}\hskip 8.53581pt\mathrm{and}\hskip 8.53581pt\mathcal{B}=\begin{bmatrix}\mathbf{B}&\mathbf{0}&\ldots&\mathbf{0}\\ \mathbf{AB}&\mathbf{B}&\ldots&\mathbf{0}\\ \vdots&\ddots&\ddots&\vdots\\ \mathbf{A}^{N-1}\mathbf{B}&\ldots&\mathbf{AB}&\mathbf{B}\end{bmatrix}\hskip 5.69054pt\mathrm{and}\hskip 8.53581pt\mathcal{Z}=\begin{bmatrix}\mathbf{1}&\mathbf{0}&\ldots&\mathbf{0}\\ \mathbf{A}&\mathbf{1}&\ldots&\mathbf{0}\\ \vdots&\ddots&\ddots&\vdots\\ \mathbf{A}^{N-1}&\ldots&\mathbf{A}&\mathbf{1}\end{bmatrix}.$

### v2

@@ Abstract
Autonomous vehicles must negotiate with pedestrians in ways that are not only safe, but also legible and socially acceptable. This paper presents an interaction-aware model predictive decision-making (IAMPDM) framework that couples a gap-acceptance-inspired pedestrian intention model with model predictive control (MPC) to coordinate intent reasoning and vehicle control in real time. The pedestrian model maps time-to-collision (TTC) to a continuous crossing-likelihood and incorporates the discounting of the pedestrian’s crossing intention to represent hesitation and potential standstills. The resulting signals then scale MPC safety costs and minimum-distance constraints. We implement IAMPDM in a projection-based, motion-tracked simulator and compare it with a rule-based intention-aware controller (RBDM) and a conservative non-interactive baseline (NIA). In a human-in-the-decision-loop study with 25 participants (one excluded from subjective analyses), intention-aware methods generally reduced scenario completion time relative to NIA in the tested scenarios, while also producing tighter surrogate safety margins (e.g., TTC- and distance-based measures). Across most metrics, IAMPDM and RBDM were not statistically distinguishable, except for TTC in one scenario. Participants’ subjective ratings tended to favor the intention-aware approaches over the non-interactive baseline. Finally, we translate these simulator observations into concrete tuning guidance.
@@ 1 Introduction
Highly automated and autonomous vehicles (AVs) are increasingly becoming a part of our daily lives Zhou et al. [2022]. Integrating these systems into society largely depends on the trust of vulnerable road users, such as cyclists and pedestrians, whose safety is essential Bouzidi and Hashemi [2023]. Incidents involving automated driving functions often make headlines, fostering public skepticism towards these technologies. Consequently, significant research efforts are directed at enhancing automated vehicles with advanced communication channels and decision-making algorithms designed to manage complex interaction scenarios, see Markkula et al. [2020a]. For example, in urban environments and city centers – where vehicles move at slower speeds, and pedestrians may cross unexpectedly – an effective human-machine interaction is critical to fostering trust in these systems. Figure 1 illustrates a typical urban scenario where a pedestrian crosses the street at an unsignalized intersection while interacting with an autonomous vehicle.
@@
Figure 1: The pedestrian crosses the street at an unsignalized intersection interacting with an autonomous vehicle. With the courtesy of version1 GmbH
In recent years, advances in machine learning and control – especially game-theoretic models and reinforcement learning (RL) – have improved automated vehicles’ ability to interact with pedestrians and cyclists by capturing the negotiation and decision dynamics that arise in mixed-traffic situations, Hu et al. [2018], Razmi Rad et al. [2020], Kampitakis et al. [2023], Kalantari et al. [2023] The performance and safety of these interaction strategies reported in the literature are commonly evaluated under two principal paradigms. The first is pure simulation - large-scale, repeatable virtual environments that allow researchers to explore many scenarios and parameter settings efficiently Luu et al. [2022], Kampitakis et al. [2023]. The second paradigm embeds humans into the evaluation loop, either through human-in-the-loop studies or controlled user experiments, which can reveal important aspects of human behavior, trust, and perceived safety that simulations alone cannot capture Tran et al. [2021], Bindschädel et al. [2021].
While both approaches are indispensable, literature-based experiments typically impose constraints on the decision space: vehicle actions (e.g., braking profiles, trajectory choices) and the range of participants’ responses are often predefined to make analysis tractable and results reproducible. This strong experimental control facilitates statistical comparison across algorithms and conditions but comes at the cost of external validity - the restricted set of behaviors and responses can fail to capture the unpredictability and heterogeneity of real-world interactions. As a result, findings derived from such controlled setups may overestimate robustness or underrepresent rare but safety-critical behaviors, limiting the applicability of conclusions when algorithms are deployed outside laboratory conditions.
Therefore, in this work, we address those limitations by proposing and validating an interaction aware model predictive decision-making (IAMPDM) approach in a simulator using a true human-in-the-decision-loop setup where participants are free to choose their actions. Simulator studies are a practical and ethically safer alternative to on-road trials for examining unconstrained human–AV negotiation Tran et al. [2021]. Building on this rationale, the present paper reports a simulator experiment that preserves participant freedom while enabling a systematic comparison of decision-making strategies. The main contributions are:
@@
    We adapt the IAMPDM for real-time experiments,
@@
    We design a novel human-in-the-decision-loop simulator study that balances naturalistic validity and experimental control.
@@
    Finally, we conduct an experiment with 25 participants, and analyze the findings to derive recommendations for real - world implementation of IAMPDM.
@@ 2 State of the Art
The upcoming sections review key conceptual frameworks, contrasts model-free and model-based approaches, and discusses validation methods for autonomous vehicle decision-making algorithms, with emphasis on interaction-aware systems.
2.1 Conceptual Frameworks for Interaction
Defining “interaction” in the context of mixed human-  automated traffic is a central challenge for the development of advanced driver-assistance and autonomous vehicle systems. According to the conceptual framework presented by Markkula et al. [2020a], interaction encompasses any situation in which the actions of one road user are shaped by the observed or anticipated behavior of another. This includes not only direct collision avoidance but also the exchange of information, either implicitly through kinematic cues (such as deceleration profiles or time-to-collision), or explicitly via signals and gestures. The framework of Markkula et al. [2020a] distinguishes between implicit communication - for example, the yielding intent communicated by gradual deceleration - and explicit communication, such as eye contact or hand gestures. The literature increasingly emphasizes the importance of both, as urban traffic negotiation is a continuous, communicative process involving prediction, negotiation, and mutual adaptation.
Recent surveys, Wang et al. [2022] and Crosato et al. [2024], catalogue the broad range of computational models developed for social interaction in AVs. These reviews cover approaches from game theory, which model strategic reasoning between agents, to frameworks grounded in social psychology, which seek to explain and simulate human-like negotiation, cooperation, and norm adherence in traffic.
@@ 2.2 Decision-Making Concepts
2.2.1 Model-Free, Learning-Based Approaches
@@
Model-free approaches do not rely on physics-based or logic-based models. Instead, they learn optimal behaviors through trial and error by directly interacting with the environment, embodying an end-to-end learning paradigm. These methods leverage environmental feedback to refine decision-making policies, often employing algorithms such as Q-learning or policy gradient methods. A comprehensive review of this topic can be found in Di and Shi [2021].
@@
In Russo et al. [2021], an RL approach is presented that addresses pedestrian collision avoidance in autonomous driving systems. The focus lies on managing unexpected pedestrian crossings while maintaining trajectory tracking performance. The Deep Deterministic Policy Gradient (DDPG) algorithm is utilized to learn continuous control actions in a high-dimensional state space. The authors discuss the reward function design, agent architecture, and environment model used for training and testing the DDPG-based agent through numerical simulations. Similar data-driven concepts are presented in Raghuram Kadali et al. [2014], Rezaee et al. [2021], Trumpp et al. [2022] and Zhang et al. [2022], demonstrating the breadth of learning-based approaches to pedestrian interaction modeling. Additional recent approaches have incorporated social value orientation to shape interaction policies, enabling negotiation styles from assertive to cooperative Crosato et al. [2022]. While effective at modeling subtle dynamics, these methods remain data-intensive and challenging to verify for safety.
@@
However, the main drawback of these model-free concepts is that they cannot easily ensure safety or provide formal guarantees on the reliability required in high-risk environments such as AV-pedestrian interactions. The black-box nature of learned policies makes verification challenging, and their performance may degrade in scenarios not adequately represented in training data. This limitation means that pure learning-based systems are not yet mature enough to be safely deployed in real traffic situations without additional safeguards.
It is expected that model-free decision-making approaches will become easier to validate and verify for traffic admission through the combination of model-free and model-based elements, as discussed in Korbmacher and Tordeux [2022] and Derajić et al. [2024]. Such hybrid architectures aim to leverage the adaptability of learning-based methods while retaining the interpretability and safety guarantees of model-based approaches.
2.2.2 Mechanistic Model-Based Approaches
In contrast, mechanistic model-based approaches commonly employ analytic criteria such as gap acceptance and time-to-collision (TTC) thresholds. While these criteria are more straightforward to implement and provide a baseline for decision-making, a growing body of research demonstrates that human crossing behavior is richer and depends on a combination of cues. For example, Tian et al. [2023] found that pedestrians rely not just on TTC, but also on the detailed parameters of vehicle deceleration when assessing whether it is safe to cross.
@@ 2.2.2 Mechanistic Model-Based Approaches
In Domeyer et al. [2022], a perceptual model is established based on the geometrical relations between drivers and pedestrians, emphasizing the implications for vehicle automation. It explores the coupling of driver and pedestrian behavior. However, only a simulation analysis is provided, with no strong indications for real-world usage. Game theoretical models are presented in Ezzati Amini et al. [2024], Skugor et al. [2020] and Pavelko et al. [2022]. In these works, the interaction between the pedestrian and vehicle is modeled as a game with two players who optimize their own objective functions through their joint actions. These models can represent the interaction between the two players, but they are not suitable for predicting their future joint actions.
@@
On the other hand, in Chen et al. [2023], a model predictive control (MPC) formulation is proposed, which can predict the future behavior of the pedestrian. In Škugor et al. [2023], the model predictive formulation is extended with a stochastic component, providing a more realistic overall behavior of the proposed algorithm. More recently, perceptually-grounded predictive models have been proposed that explicitly account for how pedestrians integrate deceleration cues and gap evolution when forming crossing decisions Tian et al. [2025a], and probabilistic vehicle–pedestrian interaction models have been developed to plan optimal implicit-communication behaviors under intent uncertainty Amann et al. [2025].
Building on these findings, perceptually plausible models have been introduced that integrate multiple sensory cues and contextual information into a continuous decision process. These methods rely on a predefined model to predict future states and make decisions. Furthermore, they can handle constraints and uncertainties in the system-design phase, leading to improved safety and reliability. More detailed overviews of model-based methods can be found in Camara et al. [2021], Gulzar et al. [2021] or Predhumeau et al. [2023].
2.3 Social Compliance and Empirical Validation
Interactions in traffic are multi-agent processes that combine perception, implicit signaling, and negotiation. Social compliance (intelligibility, predictability, comfort) is crucial for acceptance: at uncontrolled crossings AVs face a trade-off between conservative and aggressive behaviors, so decision policies must balance physical safety with legibility and social objectives Tian et al. [2025b]. VR studies and human–machine interfaces (eHMI) research (text, icons, lights, kinematics) have probed trust, workload, and communication effectiveness, and work on implicit cues (e.g., head movements) shows how pedestrians infer intent Zhou et al. [2022], Löcken et al. [2019], Yang et al. [2024]. Nevertheless, many experiments constrain participant influence on AV policies, limiting mutual adaptation and leaving few studies that directly compare different decision-making algorithms under truly interactive, human-in-the-loop conditions.
@@ 2.4 Shortcoming of the State-of-the-Art Methods
Despite impressive progress, current state-of-the-art methods face several limitations. Model-free approaches often lack interpretability and formal safety guarantees, hindering certification. Model-based approaches can become computationally heavy and may not scale to dense, multi-agent scenarios, and the decision-making and control algorithms from the state of the art are therefore often too complex for real-time implementation on general automotive hardware or hard to validate for motor vehicle registration. Moreover, many experiments rely on scripted behaviors or do not put humans fully into the decision loop, limiting ecological validity — human in the decision making loop experiments with no predefined behavior are an inevitable extension for human factors analysis. Finally, algorithms rarely consider social compliance explicitly, and there is a paucity of work on quantifying safety efficiency trade-offs. Therefore, our work addresses these challenges and provides an IAMPDM that can run in real time; furthermore, this work also provides an investigation having a human‑in‑the‑decision‑making‑loop character.
3 Improvements of the Model Predictive Decision Making Algorithm
@@ 3.1 Model Description
The IAMPDM implements an explicit motion model for the pedestrians’ motion. This is necessary to make the decision-making algorithm of the AV more suitable for city-center traffic scenarios with low driving speeds. The core idea of this IAMPDM is presented in Varga et al. [2023b] that we used for our implementation; and we validated the model using the data from Yang et al. [2019].
The proposed sigmoid-based TTC model, while simplified compared to multi-cue approaches Tian et al. [2023, 2025a], captures essential gap acceptance behaviors observed in empirical studies. Research on pedestrian crossing decisions reveals that time-to-arrival estimates—closely related to TTC—serve as primary decision variables, though these estimates are systematically distorted by approach speed and individual differences Raghuram Kadali et al. [2014]. Specifically, pedestrians provide higher TTC estimates for higher approach speeds and accept smaller gaps accordingly, with age-related variations in conservativeness. Our sigmoid function’s adjustable parameter $c$ accommodates these individual differences, allowing the model to represent cautious to aggressive crossing behaviors within a unified framework.
For the explicit motion model, it is assumed that the pedestrian’s choice of speed at the next time step can be modeled by
$\dot{y}_{\mathrm{ped}}(t)=\frac{1}{1+\exp\left(-TTC_{\mathrm{MPC}}(t)+c\right)}\cdot v_{\mathrm{ped}}^{\mathrm{ref}}.$ (1)
The impact of the parameter $c$ is illustrated in Figure 2. The $TTC$ is computed as
$TTC_{\mathrm{MPC}}(t)=\frac{x_{\mathrm{ped}}(t)-x_{\mathrm{veh}}(t)}{v_{\mathrm{veh}}(t)}-\frac{y_{\mathrm{veh}}(t)-y_{\mathrm{ped}}(t)}{v_{\mathrm{ped}}^{\mathrm{ref}}}.$ (2)
@@
The model’s validity extends beyond its alignment with gap acceptance literature. By focusing on TTC as the core interaction metric, we follow established approaches in interaction-aware planning Crosato et al. [2024], while our probabilistic formulation provides advantages over deterministic models. The sigmoid mapping between TTC and crossing probability naturally handles uncertainty in pedestrian intentions—a critical aspect often overlooked in rule-based systems. Furthermore, validation against real-world trajectory data from Yang et al. [2019] demonstrates that this simplified model adequately captures crossing patterns in urban intersections, despite not explicitly modeling all perceptual cues that pedestrians might use. Importantly, the model’s simplicity enables real-time computation within MPC frameworks while maintaining behavioral plausibility. This trade-off between completeness and computational efficiency reflects broader challenges in social compliance for AVs, where perfect behavioral modeling must be balanced against practical deployment constraints Tian et al. [2025b].
The main benefit of the proposed model (1) is that the decision layer is integrated into to control layer, thus the MPC can solve the trajectory planning and decision making tasks in one step. This can lead to a significant advantage compared to the state-of-the-art methods, where the decision-making and control layers are separated, which can lead to suboptimal solutions.
Figure 2: The output of (1) with $v_{\mathrm{ped}}^{\mathrm{ref}}=1.4\frac{m}{a}$. Each curve represents a different pedestrian behavioral profile, with higher values of $c$ indicating more cautious decision-making and lower values reflecting assertive crossing behavior.
Gap distance model (see e.g. Tian et al. [2024]) inherently assumes that the pedestrian’s decision-making is based on the current gap value and assessing the safety without considering how that gap might change in the future. On the other hand, (1) modulates the interaction dynamics between pedestrian and vehicle, which evolves over time. It can account for variables such as speed, acceleration, and changing distances in a continuous manner and used for predicting interaction dynamics.
Assuming a linear dynamics of the vehicle - which is common by the applications of autonomous vehicles - and that the pedestrian moves in the $y$ direction, the following discrete dynamic system is obtained. The state vector is defined as $\boldsymbol{x}(t)=[x_{\mathrm{veh}}(t),\dot{x}_{\mathrm{veh}}(t),y_{\mathrm{ped}}(t),\dot{y}_{\mathrm{ped}}(t)]^{T}$, where $x_{\mathrm{veh}}$ and $\dot{x}_{\mathrm{veh}}$ are the position and velocity of the vehicle in the driving direction, respectively. The position and velocity of the pedestrian in the $y$ direction are denoted by $y_{\mathrm{ped}}$ and $\dot{y}_{\mathrm{ped}}$, respectively. The discrete time step is $\Delta t$. The explicit dynamics of the system is given by
@@
$+\underbrace{\begin{bmatrix}0.5\cdot{\Delta}t^{2}\\ \Delta t\\ 0\\ 0\end{bmatrix}}_{{\mathbf{B}}}u_{\mathrm{veh}}(t)+\underbrace{\begin{bmatrix}0\\ 0\\ 0\\ \frac{v_{\mathrm{ped}}^{\mathrm{ref}}}{1+e^{-TTC_{\mathrm{MPC}}(t)+c}}\end{bmatrix}}_{\boldsymbol{z}(t)}$ (3)
where the desired acceleration of the vehicle $u_{\mathrm{veh}}(t)=a_{\mathrm{des}}$} is the system input. With the model (3.1), the prediction of the future states for $N$ steps is possible using the batch formulation, see Borrelli et al. [2017]. The matrices $\mathcal{A}$, $\mathcal{B}$ and $\mathcal{Z}$ are defined as:
$\mathcal{A}=\begin{bmatrix}\mathbf{A}\\ \mathbf{A}^{2}\\ \vdots\\ \mathbf{A}^{N}\end{bmatrix}\hskip 8.53581pt\mathrm{and}\hskip 8.53581pt\mathcal{B}=\begin{bmatrix}\mathbf{B}&\mathbf{0}&\ldots&\mathbf{0}\\ \mathbf{AB}&\mathbf{B}&\ldots&\mathbf{0}\\ \vdots&\ddots&\ddots&\vdots\\ \mathbf{A}^{N-1}\mathbf{B}&\ldots&\mathbf{AB}&\mathbf{B}\end{bmatrix}\hskip 5.69054pt\mathrm{and}$
$\hskip 8.53581pt\mathcal{Z}=\begin{bmatrix}\mathbf{1}&\mathbf{0}&\ldots&\mathbf{0}\\ \mathbf{A}&\mathbf{1}&\ldots&\mathbf{0}\\ \vdots&\ddots&\ddots&\vdots\\ \mathbf{A}^{N-1}&\ldots&\mathbf{A}&\mathbf{1}\end{bmatrix}$.